DETERMINATION OF THE STRUCTURE OF NEON CYCLOPENTANONE

WEI LIN, \underline{ANDREA J. MINEI}, ANDREW H. BROOKS, WALLACE C. PRINGLE, and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

Rotational spectra of the neon cyclopentanone van der Waals complex were studied using a pulsed-jet Fabry-Perot Fourier transform microwave spectrometer. Spectra of the normal isotopomer along with those of five ^{13}C, the ^{18}O, and the ^{22}Ne substituted isotopomers were assigned in the frequency region of 6-23 GHz. The asymmetry of the neon location within the complex allows the observation of a, b, and c-type transitions. The rotational constants and centrifugal distortion constants for the parent isotopomer were determined to be A = 2728.8119(7), B = 1736.5880(5), and C = 1440.4682(5) MHz; and Δ_{J} = 15.051(7), Δ_{JK} = -13.96(4), Δ_{K} = 49.97(5), δ_{J} = 3.603(3), and δ_{K} = 6.55(8) kHz. The coordinates of the neon in the principal axis system of C_{5}H_{8}O are a = 0.91, b = 0.78, and c = 3.26 Å, where the carbonyl bond lies along the a-axis and the oxygen position is a = 2.05 Å. Comparisons between the bonding and wide amplitude motions of the rare gases in the neon verses the argon cyclopentanone complexes will be discussed.